2 Statement of Account (11 Marks) Study the following account that Rebecca received from The Fab Store for March 2019, and
Answer #1
2)
2.2.1:
A = 1840-1720 = 120
B = 1840 - 100 = 1740
C = 2690 - 2390 = 300
D = 2382 - 2182 = 200
2.2.2:
The words "30 Days (less 5%)" means that if payment
is made within 30 days of purchase, a discount of 5%
is allowed on the invoice amount.
2.2.3:
Date 10th = receipt of payment of $100 against miscel. Purchases.
Date 19th = credit note of $450 ie. reduction in account, on account of goods return or trade discount.
Date 22nd = Interest of $17 was charged to Rebecca's account of credit stands to her account.
2.2.4:
Yes, "The FAB Store" is adequately controlling the amount of credit
provided to Rebecca because (a) regularly getting payments (b) interest
is being charged for the credit provided ('c) trade discount is given on
account through credit note.